OpenStudy (anonymous):

lim x-> infinity (1+e^3x)^1/5x Could someone help me solve this question. Thanks.

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

Answer is 1 Since 1/5x goes to zero when x goes to infinity and anything with the power zero is equal to 1

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty}(1+e^{3x})^{\frac{x}{5}}\]?

OpenStudy (anonymous):

Yes, but its 1/5x instead of x/5.

OpenStudy (anonymous):

\[\lim_{x\rightarrow \infty}(1+e^{3x})^{\frac{1}{5x}}\]?

OpenStudy (anonymous):

Yep :D

OpenStudy (jamesj):

Then the answer is not 1 nor 3/5. Evaluate the limit of the log of this expression.

OpenStudy (anonymous):

what jamesj said. the limit of the log is 3/5

OpenStudy (anonymous):

@jamesj i have a question maybe you would be willing to answer in chat

OpenStudy (anonymous):

So would I just plug infinity in all of the x?

OpenStudy (anonymous):

no the way you want to solve a limit where the variable is in the exponent is first take the log, then take the limit, then "exponentiate"

OpenStudy (anonymous):

The answer is e^(3/5)

OpenStudy (anonymous):

i.e. take the limit of \[\frac{1}{5x}\ln(1+e^{3x})\]

OpenStudy (jamesj):

yes exactly. (@sat73, don't know if you missed my comment in chat; it's there.)

OpenStudy (anonymous):

you get \[\frac{\infty}{\infty}\] and then you can use l'hopital

OpenStudy (anonymous):

oh i missed it but i will look, thanks

OpenStudy (anonymous):

Oh okay, alright thank you for your help.